# Cubic BeB2: A metastable -type conductive material from first principles

### v1

@@ I Introduction
Boron is a element that sits in a unique place in the periodic table. As a group III element, it forms III-V compound with N, P, and As, whose functional properties have been consistently under investigation for decades. In addition, compounds formed between boron and electropositive elements from Groups I and II display diverse structures and electronic behaviors, owing to the unique bonding preferences of boron and the charge-donating role of the metals [1]. An example is MgB2, one of the most extensively studied materials for high-temperature superconductivity [2, 3, 4]. In this compound, electron donation from Mg enables the formation of a graphene-like B network. While MgB2 has shown remarkable electronic properties, the relatively large size of Mg limits its ability to form denser or more compact polymorphs with B, particularly those featuring 3D covalent boron networks. In addition, metal borides have also been investigated as potential thermoelectric materials due to their ability to withstand high temperatures and relatively high values [5].
@@
Among the boride materials, the compounds between Be, which is one row above Mg in the periodic table, and B, exhibit an interesting and complex behavior. The small atomic size of Be enables the formation of a wider variety of structural motifs in Be–B compounds, favoring more compact and tightly bonded configurations. Previous studies have identified a rich structural landscape for potential Be-B binaries, ranging from molecular to crystalline compound, and spanning from B- to Be-rich compositions [6, 7, 1]. From a chemical perspective, Be-B compounds with 1:2 stoichiometry are particularly interesting members among Be-B binaries. The electronic configuration in BeB2 enables the boron atom to behave similarly to a carbon atom, leading to structures that closely resemble those formed by carbon. The similarity has been demonstrated both at the crystalline level, where at least five metastable BeB2 phases have been predicted theoretically [8], and in nanostructures, where carbon-like nanocrystals have been observed [9]. Although the BeB2 hexagonal phase, which is an MgB2 analog, has not shown a superconductivity behavior as promising as MgB2 [10, 11, 12], the 2D form (monolayer) of BeB2 has recently been studied as a potential anode material for battery applications [13, 14, 15].
@@
Among the various Be-B compounds, cubic BeB2 (-BeB2) stands out due to its distinctive crystal structure. As shown in Fig. 1, -BeB2 adopts a face-centered cubic (FCC) lattice with a Pearson symbol cF12 and belongs to the space group F4m (No. 216). The formation of this structure is made possible by the interplay between the small atomic size of Be and the valence states of both Be and B. In -BeB2, each Be atom donates two electrons to its neighboring B atoms, resulting in a four-electron valence configuration per B atom. The two B atoms in the unit cell are not equivalent due to the different distance to the Be atom, resulting in the formation of a zinc-blende-like network. Previous studies have identified -BeB2 as a metastable phase, thermodynamically stable only under high pressure conditions [8, 6, 1], posing challenges for experimental synthesis. However, its unique bonding character warrants a comprehensive ab initio investigation to explore its functional properties and underlying structural-property relationships. The diamond-like crystal structure suggests high hole mobility, but in contrast to diamond, which has very high acceptor ionization energy, the metastable phase of BeB2 is expected to have high hole concentration, even intrinsic -type behavior. Furthermore, inspired by the discovery of superconductivity in heavily doped diamond [16, 17], Si [18, 19], and SiC [20, 21], the electronic structure of -BeB2, with boron atoms acting as pseudo-carbons, presents an opportunity for further exploration of its superconducting properties.
@@ II Computational details
The electronic structure is obtained using the PBE exchange-correlational functional, the hybrid functional technique based on the Heyd-Scuseria-Ernzerhof (HSE06) [28, 29] scheme, as well as considering quasiparticle effects through many-body perturbation theory (the approximation). We evaluate the quasiparticle energies using the BerkeleyGW package [30, 31]. A BZ sampling of is used to evaluate the quasiparticle energies. Subsequently, the maximally localized Wannier function (MLWF) technique [32] is used to obtain a smooth electronic band structure. Hybrid functional calculations are performed with the VASP package [33, 34], using the projector augmented wave (PAW) treatment of core electrons [35].
To calculate direct optical properties, we solve the Bethe-Salpeter equation for optical polarization functions to take electron-hole Coulomb interactions into account [36]. A fine BZ sampling grid of is used to evaluate optical properties, with the energies and Coulomb matrix elements interpolated by considering wavefunction overlaps [30] from the coarse grid.
We use density functional perturbation theory (DFPT) [37] to compute the dynamical matrix and the electron-phonon matrix elements on a BZ grid. The electron-phonon matrix elements are subsequently interpolated on a fine grid using the MLWF technique to enable accurate carrier transport calculations. The hole mobility is determined by solving the Boltzmann transport equation, as implemented in the EPW code [38, 39, 40]. Both hole-phonon scattering and hole-ionized-impurity scattering are considered within the framework of the point charge model [41] since in -type semiconductors, the latter usually dominates at moderate to high doping levels. A fine grid of is used for transport calculations, and convergence tests confirm that increasing the grid density to results in mobility changes of less than 1%. For superconductivity calculations, we use for and -point grids, with an energy window of eV around the Fermi level. For calculating the isotropic Eliashberg spectral functions, the Dirac deltas of electrons and phonons are replaced by Gaussians of width 50 meV and 0.1 meV, respectively.
@@
and are the total energies of the defect and pristine supercells, respectively. and denote the number of defect atoms and the corresponding chemical potential. and are the Fermi level and the energy of the valence band maximum (VBM), respectively. represents the charge state of the defect. is a correction induced by considering periodic charge, and we adopt the scheme of Freysoldt et al. [42], with the additional corrections due potential alignment evaluated using the pymatgen-analysis-defects package [43].
@@ III Electronic, vibrational, and optical properties
The relaxed lattice parameter of -BeB2 is found to be Å with the PBE exchange-correlation functional, and Å with the HSE hybrid functional. The calculated BB bond length is 1.86 Å, slightly longer than the CC bond length in diamond (1.6 Å from simulation and 1.55 Å from experiments [44]). The formation enthalpy is evaluated as , yielding a value of 0.018 eV/atom, in excellent agreement with the literature value of 0.017 eV/atom [8, 6]. An early study on Be-B compounds reported -BeB2 as the ground-state structure at this composition under ambient pressure. However, it only lies on the convex hull above 80 GPa [6], due to the presence of the thermodynamically stable BeB2.75 phase at ambient conditions. Subsequent work identified a lower-energy BeB2 phase at ambient pressure that crystallizes in the orthorhombic structure (oS12, No. 63) [1, 11, 8], though this phase remains metastable with respect to the convex hull by 13 meV/atom. Consequently, -BeB2 (cF12) is a higher energy metastable phase, lying 85 meV/atom above BeB2 (oS12) phase and 98 meV above the convex hull [1] at ambient pressure. We verified that the material remains dynamically stable and no significant changes in the electronic structure are seen at a very high pressure of 80 GPa (see supplemental materials Fig. S1[45]). However, the significant energy offset suggests that bulk synthesis of -BeB2 is difficult to achieve experimentally, consistent with the experimental observation of a large uncertainty of phases in Be-B compounds [7]. Despite of the large energy barrier, the 4.297 Å lattice constant of -BeB2 closely matches those of several common substrates, including 3C-SiC (4.359 Å) [46] and MgO (4.212 Å) [47]. Additionally, the equivalent lattice constant in the (111) hexagonal plane of -BeB2 is 3.038 Å, which aligns well with the in-plane lattice constant of 4H-SiC (3.073 Å) [46]. As a result of this favorable lattice match to cubic systems, epitaxial stabilization can serve as a possible route for realizing this metastable phase in thin-film form.
@@
We further computed the electronic band structure and phonon dispersion of -BeB2, shown in Fig. 2. The electronic band gap values at different levels of theory are listed in Table 1, along with comparisons to literature reports. A comparison of the band structure from PBE and HSE exchange-correlation functionals is shown in Fig. S2 [45]. The band gap evaluated with the PBE exchange-correlation functional agrees well with previously reported values [8, 6]. Notably, the -BeB2 phase remains the only semiconducting phase among all known polytypes. An indirect band gap is found consistently using the PBE and HSE exchange-correlation functionals, as well as with single-shot based on the PBE calculations. While PBE is known to severely underestimate the gap, the band gap from the HSE exchange-correlation functional (1.52 eV) and the (1.58 eV) approximation are found to be very close. The difference between the indirect band gap, located along the direction, compared to the minimum direct gap, located at the point, is found to be approximately 0.23-0.28 eV. The density of state evaluated from the HSE exchange-correlation functional shows that the valence band maximum has the main characters of boron state. Both boron atoms contribute comparably up to 0.5 eV below the VBM, below which the contribution from the more distant B2 becomes more dominant. The conduction band minimum, on the other hand, is mainly composed of states from the Be atom and the closer B1 atom, suggesting that B1 receives more electrons from Be compared to B2. This feature is further confirmed by the calculation of Born effective charges discussed below.
@@
We evaluate the hole effective masses by performing a parabolic fits to the electronic band structure calculated with both the PBE and the HSE exchange-correlation functionals. We consider states within 5% of the VBM along the three high-symmetry directions of the BZ: (100), (110), and (111). The resulting effective masses are listed and compared to those of other common FCC semiconductors in Table 2. We find a pronounced difference between the light and heavy hole effective masses in -BeB2, more so than in materials such as SiC, diamond, and -BAs. In particular, the light holes exhibit significantly lower effective masses across all directions, suggesting the potential for efficient hole transport in this material. The trend is consistent between PBE and HSE. Quantitatively, the inclusion of the Hartree exact exchange in HSE widens the VBM, therefore increasing the effective masses in general. However, the effective masses of the light holes remain very similar in all directions, with the most pronounced impact of HSE on the heavy holes along the direction.
@@
Phonon dispersion calculations reveal no imaginary phonons, indicating that -BeB2 is dynamically stable. The high-frequency dielectric constant, calculated using DFPT, is found to be 13.99. Longitudinal-optical–transverse-optical (LO-TO) splitting at the -point is observed due to the polar nature of the BeB bonds. The computed Born effective charges are +2.2 on the Be atom, 2.1 on the B1 atom nearest to Be, and 0.1 on the more distant B2 atom. The phonon density of state indicates that the majority of the low frequency optical phonon modes are associated with the charged boron (B1) while the high frequency optical modes are associated with the neutral boron atom (B2).
@@
Given the range of the band gaps evaluated using both hybrid functionals and the approximation, -BeB2 appears to be a feasible candidate for optoelectronic applications. In principle, evaluating optical absorption across an indirect band gap requires accounting for phonon-assisted transitions. However, the difference between the minimum direct gap and the fundamental indirect gap is only 0.28 eV at the level. At finite temperature, the broadening of electronic states and spectral features is expected to render direct transitions dominant. Therefore, our analysis focuses on direct optical absorption only. The calculated optical absorption spectra for both pristine -BeB2 and heavily -type doped samples with carrier concentrations of cm-3 and cm-3 are presented in Fig. 3. In the pristine material, a weak excitonic effect is evident by comparing spectra with and without the inclusion of electron-hole Coulomb interactions. The sharp increase in absorption beyond the absorption onset suggests that the material can be an excellent absorber of visible light. In heavily -doped samples, two effects are observed due to the partially empty VBM: First, a gap opening is seen due to the shifting down of the highest occupied level, and second, a strong sub-gap absorption due to inter-valence optical transitions. An evaluation of the absorption coefficient shows that in the phonon energy range of 2-4 eV, the material exhibits an absorption coefficient exceeding cm-1 regardless of the impact of the doping level on the band gaps (see Fig. S3[45] for the calculated absorption coefficient). As a result, despite a strong band gap opening due to Burstein-Moss effect at heavy doping, the material is not a likely candidate for transparent conducting applications.

### v2

@@ I Introduction
Boron is an element that holds a unique place within the periodic table. As a group III element, it forms III-V compounds with N, P, and As, whose functional properties have been consistently under investigation for decades. In addition, compounds formed between boron and electropositive elements from Groups I and II display diverse structures and electronic behaviors, owing to the unique bonding preferences of boron and the charge-donating role of the metals [1]. An example is MgB2, one of the most extensively studied materials for high-temperature superconductivity [2, 3, 4]. In this compound, electron donation from Mg enables the formation of a graphene-like B network. While MgB2 has shown remarkable electronic properties, the relatively large size of Mg limits its ability to form more compact polymorphs with B, particularly those featuring 3D covalent boron networks. In addition, metal borides have been investigated as potential thermoelectric materials due to their ability to withstand high temperatures and relatively high values [5].
@@
Among the boride materials, the compounds between Be, which is one row above Mg in the periodic table, and B, exhibit an interesting and complex behavior. The small atomic size of Be enables the formation of a wider variety of structural motifs in the Be-B compounds, allowing for more compact and tightly bonded configurations. Previous studies have identified a rich structural landscape for potential Be-B binaries, ranging from molecular to crystalline compounds, and spanning from B- to Be-rich compositions [6, 7, 1]. From a chemical perspective, Be-B compounds with 1:2 stoichiometry are particularly interesting members among the Be-B binaries. The electronic configuration in BeB2 enables the boron atom to behave similarly to a carbon atom, leading to structures that closely resemble those formed by carbon. This similarity has been demonstrated both at the crystalline level, where at least five metastable BeB2 phases have been predicted theoretically [8], and in nanostructures, where carbon-like nanocrystals have been observed [9]. Although the hexagonal phase of BeB2, which is an MgB2 analog, has not shown a superconductivity behavior as promising as MgB2 [10, 11, 12], the 2D form (monolayer) of BeB2 has recently been studied as a potential anode material for battery applications [13, 14, 15].
@@
Among the various Be-B compounds, cubic BeB2 (-BeB2) stands out due to its distinctive crystal structure. As shown in Fig. 1, -BeB2 adopts a face-centered cubic (FCC) lattice with a Pearson symbol cF12 and belongs to the space group F4m (No. 216). The formation of this structure is made possible by the interplay between the small atomic size of Be and the valence states of both Be and B. In -BeB2, each Be atom donates two electrons to its neighboring B atoms, resulting in a four-electron valence configuration per B atom. The two B atoms in the unit cell are not equivalent due to the different distances from the Be atom, resulting in the formation of a zinc-blende-like network. Previous studies have identified -BeB2 as a metastable phase, thermodynamically stable only under high pressure conditions [8, 6, 1], posing challenges for experimental synthesis. However, its unique bonding character warrants a comprehensive ab initio investigation to explore its functional properties and the underlying structural-property relationships. The diamond-like crystal structure suggests high hole mobility, but in contrast to diamond, which has very high acceptor ionization energy, the metastable phase of BeB2 is expected to have a high hole concentration, even intrinsic -type behavior. Furthermore, inspired by the discovery of superconductivity in heavily doped diamond [16, 17], Si [18, 19], and SiC [20, 21], the electronic structure of -BeB2, with boron atoms acting as pseudo-carbons, presents an opportunity for further exploration of its superconducting properties.
@@ II Computational details
The electronic structure is obtained using the PBE exchange-correlation functional, the hybrid functional technique based on the Heyd-Scuseria-Ernzerhof (HSE06) [28, 29] scheme, as well as considering quasiparticle effects through many-body perturbation theory (the approximation). We evaluate the quasiparticle energies using the BerkeleyGW package [30, 31]. A BZ sampling of is used to evaluate the quasiparticle energies. The dielectric screening is evaluated using the random-phase approximation, and its frequency dependence is approximated with the generalized plasmon-pole model of Hybertsen and Louie[31]. The static remainder approach is used to treat the summation over empty bands[30]. A cutoff of 40 Ry is used for evaluating the screened Coulomb interaction. Convergence of the quasiparticle calculations with respect to the number of summed empty states is shown in the supplemental information (Fig.S1)[32]. Subsequently, the maximally localized Wannier function (MLWF) technique [33] is used to obtain a smooth electronic band structure. Hybrid functional calculations are performed with the VASP package [34, 35], using the projector augmented wave (PAW) treatment of core electrons [36].
To calculate direct optical properties, we solve the Bethe-Salpeter equation for optical polarization functions to take electron-hole Coulomb interactions into account [37]. A fine BZ sampling grid of is used to evaluate optical properties, with the energies and Coulomb matrix elements interpolated by considering wavefunction overlaps [30] from the coarse grid. The optical spectra of doped samples are evaluated by rigidly shifting the Fermi level due to the change in electron occupations, while neglecting effects such as partial ionization of dopants, or the modification of the band structure by the dilute dopant concentrations considered.
We use density functional perturbation theory (DFPT) [38] to compute the dynamical matrix and the electron-phonon matrix elements on a BZ grid. The electron-phonon matrix elements are subsequently interpolated on a fine grid using the MLWF technique to enable accurate carrier transport calculations. The hole mobility is determined by solving the Boltzmann transport equation, as implemented in the EPW code [39, 40, 41]. Both hole-phonon scattering and hole-ionized-impurity scattering are considered within the framework of the point charge model [42] since in -type semiconductors, the latter usually dominates at moderate to high doping levels. A fine grid of is used for transport calculations, and convergence tests confirm that increasing the grid density to results in mobility changes of less than 1%. For superconductivity calculations, we use for and -point grids, with an energy window of eV around the Fermi level. For calculating the isotropic Eliashberg spectral functions, the Dirac deltas of electrons and phonons are replaced by Gaussians of width 50 meV and 0.1 meV, respectively. The zero-point and finite-temperature renormalization of the band gap due to electron-phonon interactions is evaluated using the Wannier function perturbation theory[43] with phonon -grid, which converges the renormalization value within 0.3%. The well-behaved convergence is consistent with recent results reported for diamond[44].
@@
and are the total energies of the defect and pristine supercells, respectively. and denote the number of defect atoms and the corresponding chemical potential. and are the Fermi level and the energy of the valence band maximum (VBM), respectively. represents the charge state of the defect. is a correction induced by considering periodic charge, and we adopt the scheme of Freysoldt et al. [45], with the additional corrections due potential alignment evaluated using the pymatgen-analysis-defects package [46]. The ranges of elemental chemical potentials are determined by the formation enthalpies of -BeB2 and the two secondary phases associated with the dopants considered, LiB and BH3, evaluated using the HSE06 exchange-correlation functional.
@@ III Electronic, vibrational, and optical properties
The relaxed lattice parameter of -BeB2 is found to be Å with the PBE exchange-correlation functional, and Å with the HSE hybrid functional. The calculated BB bond length is 1.86 Å, slightly longer than the CC bond length in diamond (1.6 Å from simulation and 1.55 Å from experiment [47]). The bond length of BeB1 is 1.86 Å, and the bond length of BeB2 is 2.15 Å. The formation enthalpy is evaluated as , yielding a value of 0.018 eV/atom (0.054 eV/f.u.) with the HSE06 hybrid functional, in excellent agreement with the literature value of 0.017 eV/atom [8, 6]. An early study on Be-B compounds reported -BeB2 as the ground-state structure at this composition under ambient pressure [6]. However, it only lies on the convex hull above 80 GPa [6], due to the presence of the thermodynamically stable BeB2.75 phase at ambient conditions. Subsequent work identified a lower-energy BeB2 phase at ambient pressure that crystallizes in the orthorhombic structure (oS12, No. 63) [1, 11, 8], though this phase remains metastable with respect to the convex hull by 13 meV/atom. Consequently, -BeB2 (cF12) is a higher energy metastable phase, lying 85 meV/atom above BeB2 (oS12) phase and 98 meV/atom above the convex hull [1] at ambient pressure. We verified that the material remains dynamically stable and no significant changes in the electronic structure are seen at a very high pressure of 80 GPa (see supplemental materials Fig. S2[32]). However, the significant energy offset suggests that bulk synthesis of -BeB2 is difficult to achieve experimentally, consistent with the experimental observation of a large uncertainty of phases in Be-B compounds [7]. Despite the large energy barrier, the 4.297 Å lattice constant of -BeB2 closely matches those of several common substrates, including 3C-SiC (4.359 Å, -1.4%) [48] and MgO (4.212 Å, 1.9%) [49]. Additionally, the equivalent lattice constant in the (111) hexagonal plane of -BeB2 is 3.038 Å, which aligns well with the in-plane lattice constant of 4H-SiC (3.073 Å, -1.1%) [48]. As a result of this favorable lattice match to cubic systems, epitaxial stabilization can serve as a possible route for realizing this metastable phase in thin-film form. Importantly, metastable cubic BeB2 and the thermodynamically stable BeB2.75 polytype (which lies on the convex hull) adopt markedly different crystal structures. BeB2.75 crystallizes in a large hexagonal unit cell (consisting of 110 atoms) that is composed of a complex network of icosahedral boron clusters (, space group No. 191). This structure is incompatible with high-symmetry cubic substrates, such as MgO, 3C-SiC. Moreover, its significantly larger hexagonal lattice constant ( Å[50]) is mismatched by 5.7% compared to the hexagonal lattice constant of 4H-SiC. Therefore, BeB2.75 is less likely to nucleate coherently under epitaxial conditions on these substrates, with which cubic BeB2 is nearly lattice matched.
@@
We further computed the electronic band structure and phonon dispersion of -BeB2, shown in Fig. 2. The electronic band gap values at different levels of theory are listed in Table 1, along with comparisons to literature reports. A comparison of the band structure from PBE and HSE exchange-correlation functionals is shown in Fig. S3 [32]. The band gap evaluated with the PBE exchange-correlation functional agrees well with previously reported values [8, 6]. Notably, the -BeB2 phase remains the only semiconducting phase among all known polytypes. An indirect band gap is found consistently using the PBE and HSE exchange-correlation functionals, as well as with single-shot based on the PBE calculations. While PBE is known to severely underestimate the gap, the band gap from the HSE exchange-correlation functional (1.52 eV) and the (1.58 eV) approximation are found to be very close. The difference between the indirect band gap, located along the direction, compared to the minimum direct gap, located at the point, is found to be approximately 0.23-0.28 eV. The electron density of states evaluated from the HSE exchange-correlation functional shows that the valence band maximum is composed primarily of boron orbitals. Both boron atoms contribute comparably up to 0.5 eV below the VBM, below which the contribution from the B1 atoms (which are closer to Be) becomes more dominant. The conduction band minimum, on the other hand, is mainly composed of states from the Be atoms and the B2 atoms (the ones farther from Be). By analyzing the Born effective charges below, we provide a more detailed insight on the relationship of the bonding and the distance of the atoms.
@@
We evaluate the hole effective masses by performing parabolic fits to the electronic band structures calculated with both the PBE and the HSE exchange-correlation functionals. We consider states within 5% of the VBM along the three high-symmetry directions of the BZ: (100), (110), and (111). The resulting effective masses are listed and compared to literature values for other common cubic semiconductors in Table 2. We find a pronounced difference between the light and heavy hole effective masses in -BeB2, more so than in materials such as SiC, diamond, and -BAs. In particular, the light holes exhibit significantly lower effective masses across all directions, suggesting the potential for efficient hole transport in this material. The trend is consistent between PBE and HSE. Quantitatively, the inclusion of the Hartree exact exchange in HSE widens the bands near the VBM, therefore increasing the effective masses in general. However, the effective masses of the light holes remain similar along all directions, with the most pronounced impact of HSE on the heavy holes along the direction.
@@
Phonon dispersion calculations reveal no imaginary phonons, indicating that -BeB2 is dynamically stable. The high-frequency dielectric constant, calculated using DFPT, is found to be 13.99. Longitudinal-optical–transverse-optical (LO-TO) splitting at the -point is observed due to the polar nature of the BeB bonds, resulting in a static dielectric constant considering ionic contribution of 19.28. The computed Born effective charges are +2.2 on the Be atom, 2.1 on the B1 atom nearest to Be, and 0.1 on the more distant B2 atom. The phonon density of states indicates that the majority of the low-frequency optical phonon modes are associated with the neutral boron (B2) while the high-frequency optical modes are associated with the charged boron atom (B1), which is expected due to the stronger interaction. We further characterize the zero-point renormalization and the band gap reduction at finite temperature due to electron-phonon interactions, shown in Fig. 3. Within the PBE exchange-correlation functional, a zero-point renormalization of 189 meV is seen, which increases as a function of temperature due to stronger phonon population. At 300 K, the renormalization is increased to 232 meV.
@@
Given the range of the band gaps evaluated using both hybrid functionals and the approximation, -BeB2 appears to be a feasible candidate for optoelectronic applications. In principle, evaluating optical absorption across an indirect band gap requires accounting for phonon-assisted transitions. However, the difference between the minimum direct gap and the fundamental indirect gap is only 0.28 eV at the level. At finite temperature, the broadening of electronic states and spectral features is expected to render direct transitions dominant. Therefore, our analysis focuses on direct optical absorption only. The calculated optical absorption spectra for both pristine -BeB2 and heavily -type doped samples with carrier concentrations of cm-3 and cm-3 are presented in Fig. 4. In the pristine material, a weak excitonic effect is evident by comparing spectra with and without the inclusion of electron-hole Coulomb interactions. The sharp increase in absorption beyond the absorption onset suggests that the material can be an excellent absorber of visible light. In heavily -doped samples, two effects are observed due to the partially empty VBM: First, a gap opening is seen due to the shifting down of the highest occupied level, and second, a strong sub-gap absorption due to inter-valence optical transitions. An evaluation of the absorption coefficient shows that in the photon energy range of 2-4 eV, the material exhibits an absorption coefficient exceeding cm-1 regardless of the impact of the doping level on the band gaps (see Fig. S4[32] for the calculated absorption coefficient). As a result, despite a strong band gap opening due to Burstein-Moss effect at heavy doping, the material is not a likely candidate for transparent conducting applications.